I have a right angled triangle whose opposite and adjacent sides are given. Opposite side is 7 and the Adjacent side is 24. Find the six trigonometric ratios of the angle.
Answer by Boreal(15235)

The right triangle is a 7-24-25 one.
sin x=opp/hyp or 7/25; csc x is 25/7
cos x=adj/hyp or 24/25; sec x is 25/24
tan x=opp/adj or 7/24; ctn x is 24/7